OpenStudy (anonymous):

suppose the velocity of an automobile starting from rest is given by v =200t / 2t+8 where v is the velocity in miles per hour and t is the time in seconds. find the number of seconds when the velocity is 50 miles per hour

OpenStudy (anonymous):

reposting this because the answer is suppose to be 4 seconds can someone show or explain it

OpenStudy (anonymous):

\[50= \frac{200t}{2t +8}\]

OpenStudy (anonymous):

\[2t +8 = 4t\]

OpenStudy (akshay_budhkar):

2t+8=4t

OpenStudy (anonymous):

\[t = 4\]

OpenStudy (akshay_budhkar):

t=4

OpenStudy (anonymous):

tun-tun how about now ? did you get it

OpenStudy (anonymous):

how is it 2t+8 = 4t? i dont get it?

OpenStudy (akshay_budhkar):

divide by 50 on both sides

OpenStudy (anonymous):

Divide by 50 and multiply by 2t + 8 both sides to the expression i wrote in my first reply

OpenStudy (anonymous):

what about now?

OpenStudy (anonymous):

if i divide both sides by 50 then it will remove the 50 in the left side but it will add /50 to the right side? like = 200t / 2t+8 / 50 ?

OpenStudy (akshay_budhkar):

yes

OpenStudy (akshay_budhkar):

so its 200t/(2t+8)(50)

OpenStudy (akshay_budhkar):

which is 4t/2t+8

OpenStudy (anonymous):

ohhh now i get coz im putting 50 under 2t+8 got it got it

OpenStudy (anonymous):

hit the good answer :D

OpenStudy (anonymous):

lol ok there u go